Calculate the yield to maturity (YTM) for a one-ye
Answer #1

Answer Part 1.

Current yield = annual coupon / current price

Annual Coupon = Current Yield * Current Price = 10% * 8,000 = $800

Let YTM be i%.

8,000 = 800/(1+i) + 10,000 / (1+i)

i = 35%

Answer Part 2.

Current yield = annual coupon / current price

Annual Coupon = Current Yield * Current Price = 20% * 8,000 = $1,600

Let YTM be i%.

8,000 = 1600/(1+i) + 1600/(1+i)^2 + ... + 1600/(1+i)^20 + 10,000/(1+i)^20

8,000 = 1600 / i * (1 - (1/1+i)^20)+ 10,000/(1+i)^20

i = 20.1%

Therefore, The yield to maturity on the bond given above is greater than the YTM of a this bond.